Instant hot packs contain a solid and a pouch of water. When the pack is squeezed, the pouch breaks and the solid dissolves, increasing the temperature because of the exothermic reaction.
The following reaction is used to make a hot pack:

LiCl(s)⟶Li+(aq)+Cl−(aq)ΔH=−36.9kJ

What is the final temperature in a squeezed hot pack that contains 20.3 g of LiCl dissolved in 112 mL of water? Assume a specific heat of 4.18 J/(g⋅∘C)for the solution, an initial temperature of 25.0 ∘C, and no heat transfer between the hot pack and the environment.

Express your answer with the appropriate units.
Answer #1

20.3 g of LiCl = mass / molar mass = 20.3 g / 42.394 g/mole = 0.4788 mole.

ΔH = −36.9 kJ = - 36900 J

heat released = 36900 * 0.4788 = 17669.2 J

total mass of solution = (112 + 20.3) = 132.3 g

heat = m * s * dT

or

17669.2 = 132.3 * 4.18 * (T - 25.0)

or

T = 56.95 oC

final temperature = 56.95 oC